• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name grimm

    Log entry time 23:39:32 on July25,2005

    Entry number 149466

    keyword=HV changes: Det3 and BLumi1

    Checking the ADC range for 40uA beam current and adjusting HV
    
    Right Arm, detector 3: 
      Before: ~52k ADC, HV 1525V
      After;  ~45k ADC, HV 1515V
    
    BLumi1
     Before: ~49k ADC, HV 420V
     After:  ~45k ADC, HV 410V